# Self-Distilled RLVR

### 03 Apr 2026 v1

@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) methods such as GRPO [grpo] have become a central paradigm for training large reasoning models [deepseekr1, kimiteam2026kimik25visualagentic], where each trajectory receives only a single scalar signal determined by the outcome from the environment. On-policy distillation (OPD) [agarwal2024gkd, lu2025onpolicy] complements this by leveraging a stronger teacher model to provide dense, token-level logits as learning signals along the student’s own sampled trajectories, enriching the trajectory-level supervision to the token level and thereby achieving faster convergence. Recent work has shown that OPD from advanced teachers can match or even outperform RLVR [mimo2025flash], establishing it as an equally compelling paradigm (see Table 1 for a systematic comparison).
Despite its effectiveness, OPD relies on a separate, typically much larger teacher model, incurring substantial computational overhead. Moreover, since OPD requires computing token-level distributions over a shared vocabulary, the teacher and student models must share the same vocabulary, significantly reducing the practical scalability of this paradigm. On-policy self-distillation (OPSD) [zhao2026opsd, hbotter2026reinforcementlearningselfdistillation_sdpo] offers an appealing alternative: a single model serves as both teacher and student, where the teacher is the same model conditioned on privileged information $r$ (such as a verified reasoning trace or environmental feedback) and the student operates solely on the input query. OPSD achieves several-fold improvements in token efficiency over GRPO without requiring any external model. However, we prove this efficiency gain is fragile: as illustrated by the red curve in Figure 1(a), performance peaks early and subsequently deteriorates, accompanied by systematic privileged information leakage, where the model explicitly appeals to an invisible reference solution during inference despite never having access to it (see Figure 2 for a representative example). These phenomena raise a natural question: why does OPD work while OPSD fails?
@@
We propose RLVR with Self-Distillation (RLSD), which instantiates this principle by repurposing the teacher from a generative target to a magnitude evaluator. Specifically, the environment reward determines the direction of each token’s update (reinforcement or penalization), while the teacher’s evidence ratio modulates only the magnitude. This decoupling anchors gradient directions exclusively to the reliable environment reward, while retaining the teacher’s dense, token-level assessment for fine-grained credit discrimination across token positions. Notably, prior approaches to token-level credit assignment, such as value-function estimation in PPO [schulman2017proximalpolicyoptimizationalgorithmsppo] and various credit assignment methods [xie2025unlocking, li2026outcome], typically require training auxiliary networks or incur substantial additional overhead, yet still produce noisy estimates. In contrast, self-distillation provides a natural and virtually cost-free source of per-token credit information, requiring only a single additional forward pass. As summarized in Table 1, RLSD is the only paradigm that simultaneously achieves on-policy training, high token efficiency, rich update signals, and environment-anchored optimization, while serving as a drop-in replacement for the uniform advantage in standard GRPO without any auxiliary loss or model. Our contributions are as follows:
@@ 2 Preliminaries
GRPO. Consider a language model $\pi_{\theta}$ trained to solve reasoning tasks. Given a question $x$, the model generates a response $y=(y_{1},\dots,y_{T})$ autoregressively. In the RLVR setting, a verifier provides a binary reward $R(x,y)\in\{0,1\}$ indicating whether the response is correct. Group Relative Policy Optimization (GRPO) [grpo] samples a group of $G$ responses $\{y^{(1)},\dots,y^{(G)}\}$ from the current policy for each question $x$, and computes a sequence-level advantage for each response relative to the group:
@@
OPD and OPSD. On-Policy Distillation (OPD) [agarwal2024gkd] addresses the sparse reward problem by having the student $\pi_{\theta}$ sample its own trajectories while a separate, typically larger, teacher model $\pi_{\hat{\theta}}$ provides dense token-level supervision. While effective, maintaining a distinct teacher model throughout training introduces significant computational overhead. On-Policy Self-Distillation (OPSD) [zhao2026opsd, hbotter2026reinforcementlearningselfdistillation_sdpo] eliminates this requirement by instantiating both roles from a single model $\pi_{\theta}$: the teacher gains its informational advantage not from greater capacity, but from conditioning on privileged information $r$ (e.g., a verified reasoning trace).
@@ 3.2.2 Gradient Structure: The Mechanism of Leakage
The proof is provided in Appendix A.2. Property (i) may suggest that the deviation is innocuous on average; however, any optimizer that computes gradients on individual samples or mini-batches, such as SGD and Adam [kingma2017adammethodstochasticoptimization], is inherently path-dependent. The zero-mean perturbations in nonlinear optimization do not necessarily cancel over the course of training.
@@ 4.1 From Distribution Matching to Credit Assignment
Step 3: Clipped credit assignment. Following the design philosophy of the clipped surrogate objective in PPO [schulman2017proximalpolicyoptimizationalgorithmsppo] and GRPO, we clip the evidence weights to bound the maximum influence of any single token:
@@ 5.1 Experimental Setup
Training Data and Benchmarks. We train our models on MMFineReason-123K [lin2026mmfinereasonclosingmultimodalreasoning], a challenging subset derived from the MMFineReason-1.8M corpus via difficulty-based filtering. Specifically, inference is performed on each training sample using Qwen3-VL-4B-Thinking with four independent rollouts, and only samples where the model fails on all attempts are retained. This conservative criterion discards trivial examples and concentrates training signal on challenging problems, yielding faster convergence and more efficient use of computational resources.
@@
We evaluate on five multimodal reasoning benchmarks spanning diverse mathematical and general reasoning capabilities. MMMU [mmmu] is a massive multi-discipline benchmark covering college-level subjects across science, engineering, and humanities, requiring both perception and domain knowledge. MathVista [mathvista] assesses mathematical reasoning grounded in visual contexts. MathVision [mathvision] extends mathematical reasoning evaluation to more complex and competition-level visual problems. ZeroBench [zerobench] is a challenging benchmark specifically designed to be unsolvable by current frontier models, providing a stress test for reasoning robustness. WeMath [wemath] evaluates fine-grained mathematical problem-solving abilities with structured difficulty levels. Together, these benchmarks provide a comprehensive and rigorous assessment of both the breadth and depth of multimodal reasoning performance. For all benchmarks, we report accuracy (%) as the evaluation metric.
@@
Models and Baselines. We conduct experiments on Qwen3-VL-8B-Instruct [qwen3vl] as our base model. We compare RLSD against the following baselines. GRPO [grpo] is a standard RLVR method that estimates token-level advantages from sparse scalar outcome rewards at the sequence level, serving as our primary RL baseline. OPSD [zhao2026opsd] is an on-policy self-distillation method where a single model acts as both teacher and student, with the teacher conditioned on privileged information (e.g., verified reasoning traces) to provide dense per-token supervision. SDPO [hbotter2026reinforcementlearningselfdistillation_sdpo] extends self-distillation to the reinforcement learning with rich feedback setting, treating the current model conditioned on environment feedback as a self-teacher to derive dense credit assignment signals. GRPO+OPSD is a straightforward combination baseline that performs model updates by linearly interpolating the GRPO and OPSD objectives with a tuned weighting coefficient, aiming to simultaneously leverage sparse environment rewards and dense self-distillation signals. We also include the Base LLM (i.e., Qwen3-VL-8B-Instruct without any post-training) as a reference to quantify the overall gains from post-training.
@@
Implementation Details. We implement our method based on the VERL [verl] and EasyR1 [easyr1] frameworks. During training, the model’s maximum context length is set to 8192, with a maximum prompt length of 4096 and a maximum response length of 4096. The same prompt and response length settings are used consistently during both training and evaluation. For GRPO, GRPO+OPSD, and RLSD, the learning rate is fixed at $1\times 10^{-6}$; for OPSD and SDPO, the learning rate is set to $1\times 10^{-5}$ following their original implementations. The training batch size is set to 256, and for each prompt, we sample 8 rollouts with a sampling temperature of 1.0. For all GRPO-based methods, we use clipping thresholds of $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$, and omit both the KL penalty loss and entropy regularization loss from the objective function. The hyperparameter $\lambda$ in our RLSD is initialized at $0.5$ and linearly decayed to 0 over the first 50 training steps, and $\epsilon_{w}$ is set to $0.2$. To maintain a stable self-distillation signal, the teacher model parameters are synchronized with the student model every 10 training steps and kept frozen in between. Regarding the privileged information required by different methods: OPSD requires verified reasoning traces, for which we use reasoning trajectories distilled from Qwen3-VL-235B-A22B-Thinking and verified as correct within the MMFineReason-123K dataset; SDPO follows its original setting in the paper, using a successful previous rollout as privileged context; RLSD requires only the final ground-truth answer without any reasoning trace, making it the least demanding in terms of privileged information. All experiments are conducted on 4 compute nodes, each equipped with 8 NVIDIA H200 140GB GPUs.
@@ 5.2.1 Main Results
Table 2 presents the evaluation results across five multimodal reasoning benchmarks. RLSD achieves the highest average accuracy in both maximum length settings (56.18% at 4K and 52.78% at 2K), outperforming the Base LLM by 4.69% and GRPO by 2.32% on average under the 4K setting. Compared to GRPO, RLSD benefits from dense token-level credit assignment, yielding notable gains on challenging mathematical datasets such as MathVista (+1.9%) and MathVision (+3.91%), where fine-grained discrimination of reasoning steps matters most. RLSD also consistently outperforms the self-distillation baselines (OPSD and SDPO) in average accuracy. This empirical gap is consistent with our theoretical diagnosis in Section 3.2: while OPSD suffers from privileged information leakage due to distribution matching under information asymmetry, RLSD extracts the dense teacher signal while anchoring the update direction to the verifier reward, thereby producing more robust reasoning improvements.
@@ 6.1 Credit Assignment in RLVR
Reinforcement learning with verifiable rewards (RLVR) methods such as GRPO [grpo] have become a standard paradigm for post-training reasoning models because they optimize against automatically checkable outcomes without human preference labels [grpo, deepseekr1]. Their central limitation is that the verifier provides a sequence-level signal, so every token in a rollout inherits the same advantage regardless of whether it corresponds to a pivotal deduction or a stylistic filler. This makes long-horizon reasoning fundamentally a credit assignment problem. A common response is to introduce process reward models (PRMs) or step-level value estimators that score intermediate reasoning steps [lightman2024lets, wang2024math, luo2024improve, chen2024step, zhang2024generative, cui2025process]. Some of these methods rely on costly human step annotations, while others replace them with automated or implicit supervision; in either case, they require auxiliary reward modeling and extra computation beyond the base policy.
More recent work seeks fine-grained credit assignment while staying strictly within a verifier-only RLVR pipeline. These methods reshape token-level updates using model-internal proxies such as entropy, uncertainty, key-token statistics, attention dynamics, or outcome sensitivity [cheng2026reasoning, wangbeyond, chen2025seed, sun2025ktaemodelfreealgorithmkeytokens, xie2025unlocking, li2025attention, chen2025beyond, li2026outcome]. While they show that uniform sequence-level advantages are unnecessarily coarse, they estimate token importance entirely from intrinsic heuristics. Our method is complementary in motivation yet fundamentally different in mechanism: rather than relying on heuristic proxies, we extract rigorous token-level assessments via self-distillation under privileged context, while keeping the update direction anchored to the verifier reward.
@@ 6.2 On-Policy Distillation
Distillation has also been explored as an alternative source of dense token-level supervision for reasoning models. In on-policy distillation (OPD), a stronger teacher evaluates the student’s own rollouts and provides token-level targets along on-policy trajectories [agarwal2024gkd, lu2025onpolicy]; recent reports suggest that strong teacher-guided distillation can even rival or complement RL-based post-training [mimo2025flash]. However, standard OPD requires maintaining an external, typically larger teacher throughout training. To eliminate this overhead, recent self-distillation methods let the same model act as both student and teacher, with the teacher conditioned on privileged information such as reference solutions or verifier feedback [zhao2026opsd, hbotter2026reinforcementlearningselfdistillation_sdpo]. Closely related on-policy distillation variants have also been explored for continual learning from demonstrations, context internalization, and reasoning compression, where the teacher is conditioned on demonstrations, auxiliary context, or conciseness instructions [shenfeld2026selfdistillation, ye2026policy, sang2026policy].

### 08 Apr 2026 v2

@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) methods such as GRPO [1] have become a central paradigm for training large reasoning models [2, 3], where each trajectory receives only a single scalar signal determined by the outcome from the environment. On-policy distillation (OPD) [4, 5] complements this by leveraging a stronger teacher model to provide dense, token-level logits as learning signals along the student’s own sampled trajectories, enriching the trajectory-level supervision to the token level and thereby achieving faster convergence. Recent work has shown that OPD from advanced teachers can match or even outperform RLVR [6], establishing it as an equally compelling paradigm (see Table 1 for a systematic comparison).
Despite its effectiveness, OPD relies on a separate, typically much larger teacher model, incurring substantial computational overhead. Moreover, since OPD requires computing token-level distributions over a shared vocabulary, the teacher and student models must share the same vocabulary, significantly reducing the practical scalability of this paradigm. On-policy self-distillation (OPSD) [7, 8] offers an appealing alternative: a single model serves as both teacher and student, where the teacher is the same model conditioned on privileged information $r$ (such as a verified reasoning trace or environmental feedback) and the student operates solely on the input query. OPSD achieves several-fold improvements in token efficiency over GRPO without requiring any external model. However, we prove this efficiency gain is fragile: as illustrated by the red curve in Figure 1(a), performance peaks early and subsequently deteriorates, accompanied by systematic privileged information leakage, where the model explicitly appeals to an invisible reference solution during inference despite never having access to it (see Figure 2 for a representative example). These phenomena raise a natural question: why does OPD work while OPSD fails?
@@
We propose RLVR with Self-Distillation (RLSD), which instantiates this principle by repurposing the teacher from a generative target to a magnitude evaluator. Specifically, the environment reward determines the direction of each token’s update (reinforcement or penalization), while the teacher’s evidence ratio modulates only the magnitude. This decoupling anchors gradient directions exclusively to the reliable environment reward, while retaining the teacher’s dense, token-level assessment for fine-grained credit discrimination across token positions. Notably, prior approaches to token-level credit assignment, such as value-function estimation in PPO [9] and various credit assignment methods [10, 11], typically require training auxiliary networks or incur substantial additional overhead, yet still produce noisy estimates. In contrast, self-distillation provides a natural and virtually cost-free source of per-token credit information, requiring only a single additional forward pass. As summarized in Table 1, RLSD is the only paradigm that simultaneously achieves on-policy training, high token efficiency, rich update signals, and environment-anchored optimization, while serving as a drop-in replacement for the uniform advantage in standard GRPO without any auxiliary loss or model. Our contributions are as follows:
@@ 2 Preliminaries
GRPO. Consider a language model $\pi_{\theta}$ trained to solve reasoning tasks. Given a question $x$, the model generates a response $y=(y_{1},\dots,y_{T})$ autoregressively. In the RLVR setting, a verifier provides a binary reward $R(x,y)\in\{0,1\}$ indicating whether the response is correct. Group Relative Policy Optimization (GRPO) [1] samples a group of $G$ responses $\{y^{(1)},\dots,y^{(G)}\}$ from the current policy for each question $x$, and computes a sequence-level advantage for each response relative to the group:
@@
OPD and OPSD. On-Policy Distillation (OPD) [4] addresses the sparse reward problem by having the student $\pi_{\theta}$ sample its own trajectories while a separate, typically larger, teacher model $\pi_{\hat{\theta}}$ provides dense token-level supervision. While effective, maintaining a distinct teacher model throughout training introduces significant computational overhead. On-Policy Self-Distillation (OPSD) [7, 8] eliminates this requirement by instantiating both roles from a single model $\pi_{\theta}$: the teacher gains its informational advantage not from greater capacity, but from conditioning on privileged information $r$ (e.g., a verified reasoning trace).
@@ 3.2.2 Gradient Structure: The Mechanism of Leakage
The proof is provided in Appendix A.2. Property (i) may suggest that the deviation is innocuous on average; however, any optimizer that computes gradients on individual samples or mini-batches, such as SGD and Adam [12], is inherently path-dependent. The zero-mean perturbations in nonlinear optimization do not necessarily cancel over the course of training.
@@ 4.1 From Distribution Matching to Credit Assignment
Step 3: Clipped credit assignment. Following the design philosophy of the clipped surrogate objective in PPO [9] and GRPO, we clip the evidence weights to bound the maximum influence of any single token:
@@ 5.1 Experimental Setup
Training Data and Benchmarks. We train our models on MMFineReason-123K [13], a challenging subset derived from the MMFineReason-1.8M corpus via difficulty-based filtering. Specifically, inference is performed on each training sample using Qwen3-VL-4B-Thinking with four independent rollouts, and only samples where the model fails on all attempts are retained. This conservative criterion discards trivial examples and concentrates training signal on challenging problems, yielding faster convergence and more efficient use of computational resources.
@@
We evaluate on five multimodal reasoning benchmarks spanning diverse mathematical and general reasoning capabilities. MMMU [14] is a massive multi-discipline benchmark covering college-level subjects across science, engineering, and humanities, requiring both perception and domain knowledge. MathVista [15] assesses mathematical reasoning grounded in visual contexts. MathVision [16] extends mathematical reasoning evaluation to more complex and competition-level visual problems. ZeroBench [17] is a challenging benchmark specifically designed to be unsolvable by current frontier models, providing a stress test for reasoning robustness. WeMath [18] evaluates fine-grained mathematical problem-solving abilities with structured difficulty levels. Together, these benchmarks provide a comprehensive and rigorous assessment of both the breadth and depth of multimodal reasoning performance. For all benchmarks, we report accuracy (%) as the evaluation metric.
@@
Models and Baselines. We conduct experiments on Qwen3-VL-8B-Instruct [19] as our base model. We compare RLSD against the following baselines. GRPO [1] is a standard RLVR method that estimates token-level advantages from sparse scalar outcome rewards at the sequence level, serving as our primary RL baseline. OPSD [7] is an on-policy self-distillation method where a single model acts as both teacher and student, with the teacher conditioned on privileged information (e.g., verified reasoning traces) to provide dense per-token supervision. SDPO [8] extends self-distillation to the reinforcement learning with rich feedback setting, treating the current model conditioned on environment feedback as a self-teacher to derive dense credit assignment signals. GRPO+OPSD is a straightforward combination baseline inspired by the MOPD approach proposed in MIMO-v2-Flash [6], which jointly optimizes reinforcement learning and distillation objectives. Specifically, we combine the GRPO loss with a distillation KL divergence loss via linear interpolation with a tuned weighting coefficient. Unlike MOPD, which relies on a separately trained expert model for distillation, we adopt the OPSD formulation by using the model’s own rollout outputs as the distillation target, thereby eliminating the need for an external expert while still leveraging dense self-distillation signals alongside sparse environment rewards. We also include the Base LLM (i.e., Qwen3-VL-8B-Instruct without any post-training) as a reference to quantify the overall gains from post-training.
@@
Implementation Details. We implement our method based on the VERL [20] and EasyR1 [21] frameworks. During training, the model’s maximum context length is set to 8192, with a maximum prompt length of 4096 and a maximum response length of 4096. The same prompt and response length settings are used consistently during both training and evaluation. For GRPO, GRPO+OPSD, and RLSD, the learning rate is fixed at $1\times 10^{-6}$; for OPSD and SDPO, the learning rate is set to $1\times 10^{-5}$ following their original implementations. The training batch size is set to 256, and for each prompt, we sample 8 rollouts with a sampling temperature of 1.0. For all GRPO-based methods, we use clipping thresholds of $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$, and omit both the KL penalty loss and entropy regularization loss from the objective function. The hyperparameter $\lambda$ in our RLSD is initialized at $0.5$ and linearly decayed to 0 over the first 50 training steps, and $\epsilon_{w}$ is set to $0.2$. To maintain a stable self-distillation signal, the teacher model parameters are synchronized with the student model every 10 training steps and kept frozen in between. Regarding the privileged information required by different methods: OPSD requires verified reasoning traces, for which we use reasoning trajectories distilled from Qwen3-VL-235B-A22B-Thinking and verified as correct within the MMFineReason-123K dataset; SDPO follows its original setting in the paper, using a successful previous rollout as privileged context; RLSD requires only the final ground-truth answer without any reasoning trace, making it the least demanding in terms of privileged information. All experiments are conducted on 4 compute nodes, each equipped with 8 NVIDIA H200 140GB GPUs.
@@ 5.2.1 Main Results
Table 2 presents the evaluation results across five multimodal reasoning benchmarks. RLSD achieves the highest average accuracy, outperforming the Base LLM by 4.69% and GRPO by 2.32% on average under the 4K setting. Compared to GRPO, RLSD benefits from dense token-level credit assignment, yielding notable gains on challenging mathematical datasets such as MathVista (+1.9%) and MathVision (+3.91%), where fine-grained discrimination of reasoning steps matters most. RLSD also consistently outperforms the self-distillation baselines (OPSD and SDPO) in average accuracy. This empirical gap is consistent with our theoretical diagnosis in Section 3.2: while OPSD suffers from privileged information leakage due to distribution matching under information asymmetry, RLSD extracts the dense teacher signal while anchoring the update direction to the verifier reward, thereby producing more robust reasoning improvements. Moreover, RLSD also outperforms the additive fusion methods GRPO+OPSD by a margin of 3.27 points. Linearly combining bounded rewards with unbounded, high-variance kl losses causes severe scale mismatch, forcing a suboptimal trade-off that often destabilizes training. In contrast, RLSD effectively circumvents these pitfalls through multiplicative modulation. By exponentiating the teacher’s signal into a dynamically bounded relative multiplier, RLSD mathematically guarantees strict sign preservation.
@@ 6.1 Credit Assignment in RLVR
Reinforcement learning with verifiable rewards (RLVR) methods such as GRPO [1] have become a standard paradigm for post-training reasoning models because they optimize against automatically checkable outcomes without human preference labels [1, 2]. Their central limitation is that the verifier provides a sequence-level signal, so every token in a rollout inherits the same advantage regardless of whether it corresponds to a pivotal deduction or a stylistic filler. This makes long-horizon reasoning fundamentally a credit assignment problem. A common response is to introduce process reward models (PRMs) or step-level value estimators that score intermediate reasoning steps [22, 23, 24, 25, 26, 27, 28, 29, 30]. Some of these methods rely on costly human step annotations, while others replace them with automated or implicit supervision; in either case, they require auxiliary reward modeling and extra computation beyond the base policy.
More recent work seeks fine-grained credit assignment while staying strictly within a verifier-only RLVR pipeline. These methods reshape token-level updates using model-internal proxies such as entropy, uncertainty, key-token statistics, attention dynamics, or outcome sensitivity [31, 32, 33, 34, 10, 35, 36, 11]. While they show that uniform sequence-level advantages are unnecessarily coarse, they estimate token importance entirely from intrinsic heuristics. Our method is complementary in motivation yet fundamentally different in mechanism: rather than relying on heuristic proxies, we extract rigorous token-level assessments via self-distillation under privileged context, while keeping the update direction anchored to the verifier reward.
@@ 6.2 On-Policy Distillation
Distillation has also been explored as an alternative source of dense token-level supervision for reasoning models. In on-policy distillation (OPD), a stronger teacher evaluates the student’s own rollouts and provides token-level targets along on-policy trajectories [4, 5, 37]; recent reports suggest that strong teacher-guided distillation can even rival or complement RL-based post-training [6]. However, standard OPD requires maintaining an external, typically larger teacher throughout training. To eliminate this overhead, recent self-distillation methods let the same model act as both student and teacher, with the teacher conditioned on privileged information such as reference solutions or verifier feedback [7, 8]. Closely related on-policy distillation variants have also been explored for continual learning from demonstrations, context internalization, and reasoning compression, where the teacher is conditioned on demonstrations, auxiliary context, or conciseness instructions [38, 39, 40, 41]. A recent concurrent work, TRRD [42], also identifies the conflict between additive KL penalties and reward maximization, proposing to inject teacher probabilities into the policy importance ratio. However, TRRD operates on the trust region anchor while RLSD directly modulates the advantage magnitude via a Bayesian evidence ratio.